In 1992 the cost of stamps went up $.15 to $.18. What is the percent increase?
Answer by stanbon(75887)

In 1992 the cost of stamps went up $.15 to $.18. What is the percent increase?
-----------------------------------------
Quantity Increase: 0.18 - 0.15 = 3 cents
Percent increase: 3/15 = 1/5 = 20%
-----------------------------
Cheers,
Stan H.